OpenStudy (anonymous):

Suppose that people's heights (in centimeters) are normally distributed, with a mean of 170 and a standard deviation of 5. We find the heights of 60 people. (You may need to use the standard normal distribution table. Round your answers to the nearest whole number.) How many would you expect to be between 170 and 175 cm tall?

OpenStudy (anonymous):

It's not to difficult if you know the basics for example how many would you expect to exceed 170 cm tall would be (b) P(X > 170) = P(Z > (170 - 175)/6) = P(Z > -5/6) = 0.79767 0.79767 x 60 = 47 people (round down) Source: http://answers.yahoo.com/question/index?qid=20101110100828AAiO8Uj

OpenStudy (anonymous):

does this help because I'm not going to simply give you the answer I'm only here to help you understand the question

OpenStudy (anonymous):

@lucascoco08